OpenStudy (anonymous):

Anyone know how to do this? Given the function f(x) = 2(x + 8), find x if f(x) = 12. 2 40 −2 14

OpenStudy (freckles):

replace f(x) with 12 and solve for x

OpenStudy (tkhunny):

Have you considered substituting and solving. 12 = 2(x+8) Go!

OpenStudy (anonymous):

@freckles ty again lolol

OpenStudy (anonymous):

x=-2 @tkhunny

OpenStudy (tkhunny):

Is that correct?

OpenStudy (freckles):

you can check it @redskychicken by asking yourself is the following true: 12=2(-2+8)

OpenStudy (anonymous):

12=2(-2+8) 12=2(6) 12=12

OpenStudy (freckles):

12=12 is a totally true equation

OpenStudy (freckles):

so you did super awesome @redskychicken